OpenStudy (amber15):

What is the perimeter of a computer graphic with sides measuring? 2√18, 2√32, 2√50, and 3√12 centimeters The perimeter is _______ centimeters.

OpenStudy (michele_laino):

here we have to compute this sum: \[\begin{gathered} perimeter = 2\sqrt {18} + 2\sqrt {32} + 2\sqrt {50} + 3\sqrt {12} = \hfill \\ = 6\sqrt 2 + 8\sqrt 2 + 10\sqrt 2 + 6\sqrt 3 = ...? \hfill \\ \end{gathered} \]

OpenStudy (amber15):

okay

OpenStudy (michele_laino):

so, simplifying that expression, we get: \[perimeter = 24\sqrt 2 + 6\sqrt 3 = ...?\]

OpenStudy (amber15):

ok

OpenStudy (michele_laino):

now, you have to use a calculator, in order to evaluate that numeric expression. What do you get?

OpenStudy (amber15):

should it be in decimal

OpenStudy (michele_laino):

yes! We have: \[\begin{gathered} \sqrt 2 \cong {\text{1}}{\text{,414}} \hfill \\ \sqrt 3 \cong {\text{1}}{\text{,732}} \hfill \\ \end{gathered} \]

OpenStudy (amber15):

okay

OpenStudy (amber15):

im lost doing these problems

OpenStudy (michele_laino):

so, substituting, we get: \[perimeter = 24\sqrt 2 + 6\sqrt 3 = \left( {24 \times {\text{1}}{\text{.414}}} \right){\text{ + }}\left( {{\text{6}} \times {\text{1}}{\text{.732}}} \right) = ...?\]

OpenStudy (michele_laino):

please continue

OpenStudy (amber15):

okay so you want me add 24 x 1414 + 6 x 1732

OpenStudy (michele_laino):

not precisely. I meant this: 24 x 1.414 + 6 x 1.732=...?

OpenStudy (michele_laino):

there 1.414 stands for (1414/1000), and: 1.732 stands for (1732/1000)

OpenStudy (amber15):

ok

OpenStudy (amber15):

so you want me times 24 x 1.414 + 6 x 1.732

OpenStudy (michele_laino):

ok!

OpenStudy (amber15):

44.328

OpenStudy (michele_laino):

that's right!

OpenStudy (amber15):

awesome i like the way you explained this

OpenStudy (michele_laino):

thanks!

OpenStudy (amber15):

i will be back i got like 5 problems im stuck on but did 50 on my own and got a good grade but gotta get these last done

OpenStudy (amber15):

i wanted to ask you how would i make that to a radical

OpenStudy (michele_laino):

it is simple, as I wrote, before, the perimeter is: \[perimeter = 2\sqrt {18} + 2\sqrt {32} + 2\sqrt {50} + 3\sqrt {12} \]

OpenStudy (amber15):

ok

OpenStudy (michele_laino):

now we can write this: \[\begin{gathered} 2\sqrt {32} = 2\sqrt {{4^2} \times 2} = 2 \times 4\sqrt 2 = 8\sqrt 2 \hfill \\ 2\sqrt {50} = 2\sqrt {{5^2} \times 2} = 2 \times 5\sqrt 2 = 10\sqrt 2 \hfill \\ 3\sqrt {12} = 3\sqrt {{2^2} \times 3} = 3 \times 2\sqrt 3 = 6\sqrt 3 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

furthermore, we have: \[2\sqrt {18} = 2\sqrt {{3^2} \times 2} = 2 \times 3\sqrt 2 = 6\sqrt 2 \] So substituting into the last above expression, we get: \[perimeter = 6\sqrt 2 + 8\sqrt 2 + 10\sqrt 2 + 6\sqrt 3 \]

OpenStudy (michele_laino):

finally, if we compute the sum amongsimilar radicals, we get: \[perimeter = 24\sqrt 2 + 6\sqrt 3 \]

OpenStudy (amber15):

oh okay

OpenStudy (amber15):

reason i asked was my problem wanted exact answer and ratical if needed

OpenStudy (michele_laino):

ok!

OpenStudy (amber15):

Thank you

OpenStudy (michele_laino):

Thank you! :)